# A Survey on Sequential Recommendation

### v1

@@ 4.2.1 Recurrent Neural Networks
Compared with traditional SR models, RNNs-based SR models not only can capture short-term sequential dependencies, but also can capture long-term sequential dependencies and non-linear dynamics. In recent years, some RNNs-based SR models have been proposed. GRU4Rec [24] applies RNNs to capture long-term dependencies in session-based data. Based on GRU4Rec, GRU4Rec+ [30] proposes a novel negative item sampling and a novel ranking loss function. Some RNNs-based SR models are designed to capture each sequence independently, thus ignoring the global structure of all sequences. To solve this problem, KrNN-P [31] incorporates neighbor sequences into RNNs to capture global relationships while refining local ones. To consider the dependencies among users’ sequences, MrRNN [32] introduces manifold regularization into RNNs based on the multi-facets of collaborative filtering. However, these RNN-based models have some shortcomings. For example, they cannot avoid vanishing gradients while handling ultra-long interaction sequences.
@@ 4.2.2 Convolutional Neural Networks
In recent years, several CNNs-based SR models have been proposed. DACNN [33] leverages CNNs to capture local sequential features while considering item sequences. Caser [34] leverages convolutional filters to capture both general preferences and sequential patterns. CosRec [35] leverages 2D convolutional neural networks to model complex interaction relationships. However, CNN-based SR models have certain shortcomings. For example, CNN-based SR models might have a large number of convolutional filters, which can result in excessive computational resource consumption.
@@ 4.2.3 Transformer
Transformer architecture has two types: single-direction Transformer (i.e., uni-Transformer) and double-direction Transformer (i.e., bi-Transformer). SASRec [1] is currently the most popular SR model, which leverages the uni-Transformer. SASRec can predict the next interacted items based on historical interacted items. BERT4Rec [36] leverages bi-Transformer, which can predict the next interacted items according to historical interacted items and future interacted items. ALBERT4Rec and DeBERTa4Rec [37] demonstrate superior performance and training efficiency compared to BERT4Rec [36]. ADT [38] learns disentangled diverse interests using a Transformer-based encoder and decoder. However, the large weights of attention mechanisms might not be accurate. Additionally, position encoding and noise inputs might negatively impact the recommendation performance. Therefore, AC-TSR [39] leverages contrastive learning to achieve noise reduction. DT4IER [40] leverages decision Transformer to solve user retention challenges. Due to the uncertainty of users’ sequential behaviors, STOSA [41] leverages stochastic Gaussian distribution to embed items and introduces a novel Wasserstein self-attention module to capture item-item relationships. MFGAN [42] uses a Transformer-based generator to recommend the next possible items and multiple discriminators to evaluate generated sub-sequences. SSE-PT [43] introduces user embeddings to realize the personalization of Transformer. STRec [44] leverages sparse Transformer to focus on the most relevant interactions while predicting target items. Meanwhile, STRec [44] replaces the self-attention mechanism with a cross-attention mechanism. MoSeR [45] captures motifs hidden in behavior sequences to model micro-structure features. PHGR [46] leverages a novel hyperbolic inner product operator to enable global (all user-item interactions) and local (each user’s interactions) graph representation learning in Poincaré ball. CBiT [47] leverages contrastive learning and bi-Transformer to learn more fine-grained preferences. DCRec [48] leverages cross-view contrastive learning to capture correlations among different users’ interaction sequences. DFAR [49] leverages a dual-interest disentangling layer to disentangle positive and negative interests for learning users’ transition patterns better. DEPS [50] leverages both users’ and items’ interaction sequences as the inputs to two Transformers for learning preferences better. MrTransformer [51] leverages a preference separation module to learn both users’ common and unique preferences. CRR-Transformer [52] applies a pre-trained transformer model to an online RL algorithm. Due to the uncertainty of interactions, DT4SR [53] leverages an elliptical Gaussian distribution to describe items and adopts Wasserstein distance to measure the similarity among distributions. Transformers4Rec [54] applies Transformer-based architectures for recommendation, such as GPT-2, XLNET, and Transformer-XL. PAUP [55] captures both short- and long-term patterns through a progressive attention distribution mechanism. In this model, long behavior sequences are segmented into a series of sub-sequences using a down-sampling convolution module. SAMRec [56] enhances the recommendation performance of Transformer from the perspective of loss geometry.
@@ 4.2.4 Diffusion Model
In recent years, several SR models based on diffusion model have been proposed. DiffuASR [57] leverages a diffusion-based framework to generate high-quality interactions guaranteeing the similarity between generated and original sequences. DiffuRec [58] is the first work that applies the diffusion model to sequential recommendation. By corrupting and reconstructing target item embeddings, DiffuRec [58] injects uncertainty during the recommendation process. Diff4Rec [59] leverages the diffusion model to augment data by adopting a curriculum scheduling strategy, including interaction augmentation and objective augmentation. Diffusion-based SR models have some weaknesses. Compared with Transformer-based SR models, diffusion-based SR models require more computational resources and training time.
@@ 4.2.5 Graph Neural Networks
GNNs-based SR models can learn complex high-order interactions and capture collaborative information. Therefore, GNNs-based SR models have received widespread attention. AutoSeqRec [60] leverages an encoder and three decoders to reconstruct user-item interaction and transition matrices. CD-GES-SASRec [61] leverages a causal graph to distinguish causal and non-causal transitions. DGSR [62] constructs a user-item interaction graph to capture personal preferences and an item-item transition graph to model transitional relationships between adjacent items. Meanwhile, DGSR [62] injects high-order connections into graphs. MSSG [63] transforms items within a sequence into a star graph. Meanwhile, MSSG introduces an additional internal node to capture global information. MSSG overcomes the over-smoothing issue and realizes a linear time complexity. However, GNNs-based SR models might not fully capture sequential patterns and can be computationally expensive to train.
@@ 4.2.6 Multi-Layer Perceptions
Compared with Transformer, leveraging MLPs to calculate is more efficient (i.e., a linear time complexity). In recent years, some MLPs-based SR models have been proposed. FMLP-Rec [64] incorporates learnable filters into MLPs architecture for reducing noise. TRIMLP [65] disables triangle neurons in MLPs to account for chronological order in users’ interaction sequences. MLM4Rec [66] learns users’ global and local preferences by incorporating convolution operators into MLPs. However, the MLPs-based SR models struggle to capture complex sequential patterns.
@@ 4.4 Hybrid Models
Some SR models demonstrate significant improvements in recommendation performance by combining different techniques. For example, combining Transformer and GNNs can achieve better performance. As shown in Table 2, each technique offers distinct advantages. Therefore, the combined models can combine the strengths of different techniques. The combined models can be roughly categorized into three classes (i.e., CNNs + Transformer, GNNs + Transformer and other combined types). Integrating various neural network architectures enables SR models to achieve enhanced recommendation performance. Several commonly used combined models are depicted in Figure 6. As depicted in Figure 6, there are four kinds of combined types. Two different techniques can be leveraged in parallel. For example, CNNs and Transformer can extract short-term dependencies and long-term dependencies from user interaction sequences, respectively. And then the two kinds of dependencies are fused together. Alternatively, the two techniques can be used in a cascaded manner. For example, GNNs can be denoted as a feature extractor for extracting features, and the extracted features can then be fed into Transformer.
@@ 4.4.1 Combination of CNNs and Transformer
Transformer and CNNs exhibit unique abilities to capture different types of dependencies. Transformer has a good ability to capture long-term dependencies. CNNs are effective at modeling short-term dependencies and multiple-faced preferences. CNNs and Transformer can be combined in a parallel manner, which is formalized as follows:
The final preferences can be fed into a prediction module for recommending personalized items for user . Meanwhile, CNNs and Transformer can be integrated in a cascaded manner, which is formalized as follows:
@@ 4.4.2 Combination of GNNs and Transformer
GNNs and Transformer excel at modeling graph structure and sequences, respectively. Transformer can capture temporal patterns and long-term dependencies. Meanwhile, GNNs can capture structural information and users’ long-term preferences. Therefore, combining GNNs and Transformer enables SR models to achieve better performance. GNNs and Transformer can be combined in a parallel manner, which is formalized as follows:
In the meantime, GNNs and Transformer can be integrated in a cascade manner, which is formalized as follows:
@@ 4.4.3 Other Combinations
RANN [81] leverages the strengths of RNNs and the self-attention mechanism in Transformer to capture user preferences from different views. RCNN [82] combines the advantages of RNNs and CNNs, in which RNNs can capture long-term dependencies and CNNs can capture short-term sequential patterns. NISRec [83] captures and aggregates the long-term and short-term intentions of neighbor users. Then NISRec [83] fuses these intentions for recommendation. Mixture networks can leverage multiple experts to learn diverse interests. For example, M3 [84] leverages a mixture of experts, where each expert specializes in capturing a specific temporal range. These experts are dynamically combined through a learned gating mechanism. Therefore, M3 [84] can capture diverse behavior patterns effectively.
By combining different techniques, SR models can achieve better performance. However, these models have several drawbacks. Firstly, they are more complex and computationally expensive. Secondly, different techniques might capture conflicting preferences. For example, Transformer and GNNs capture distinct preferences, which may conflict. Thirdly, different techniques converge at different rates.
@@ 5 SR with Side Information
In this section, our survey first discusses SR models combining IDs and one of the general features (i.e., categorical features and numerical features). Then, our survey introduces SR models combining IDs and reviews. Next, our survey shows SR models combining IDs and one of the graph-structure features (i.e., knowledge graph and social networks). Finally, our survey shows SR models combining IDs and multiple feature types.
@@
MLPs-based SR models | Computational efficiency, a linear time and space complexity, simple model architecture | Limited performance on ultra-long sequences, weak in capturing sequential dependencies |
@@
TCNs-based SR models | Can preserve sequence order during convolution process and conduct efficient sequence modeling | Are difficult to capture long-term sequential dependencies, high complexity for hyper-parameters searching |
@@ 5.1 One Categorical or Numerical Feature
Many SR models combine either a categorical or a numerical feature with item IDs to recommend items for users. Some specific methods are shown in Figure 7. From Figure 7, we can know that these methods can be grouped into four main types. The first method leverages two backbones to learn two kinds of preferences from one sequence containing item IDs and the other sequence containing feature , respectively. These preferences are then fused using methods such as contrastive learning, concatenation, addition, gate addition, or cross-attention. The fusion process is formalized as follows:
The second method leverages the value of feature to split the original sequence containing item IDs into sub-sequences. For example, some SR models split hybrid interaction sequences into separate interaction sequences in each domain [85, 86]. Different preferences are then extracted from these sub-sequences and fused together to predict personalized items. The fusion process is formalized below:
It is noted that the fusion usually exhibits a certain directionality, where user preferences are transferred from a source domain to a target domain [87]. The third method leverages simple techniques to combine item embeddings and feature embeddings. Then, the fused embeddings are fed to a backbone model for recommendation. For example, adding them together is to obtain the input of a backbone model. The fusion process is formalized as follows:
The last method integrates item embeddings with feature embeddings in a backbone model. For example, leveraging feature embeddings changes the weights in attention matrices. The fusion process is formalized as follows:
Though all four kinds of methods can obtain final user ’s preferences, they vary in terms of complexity and efficiency [87, 88]. Meanwhile, the predicted goal might contain the feature of the next interacted item [89, 25].
@@ 5.1.1 Behaviors
Different behaviors show users’ different preferences. Click behaviors usually reflect users’ short-term preferences. Favorite behaviors usually reflect users’ middle-term preferences. Purchase behaviors usually reflect users’ long-term preferences. Meanwhile, dependencies exist among these different behaviors. MB-STR [90] injects users’ behaviors into Transformer to capture multi-behavior heterogeneous dependencies. PBAT [91] models users’ personalized behavior patterns and behavioral collaborations by a personalized behavior pattern generator and a behavior-aware collaboration extractor, respectively. GHTID [92] leverages a global heterogeneous graph (constructed from all interaction sequences) and local heterogeneous graphs (constructed from each interaction sequence) to explicitly learn heterogeneous item transitions.
@@ 5.1.4 Domain
Domain information plays a crucial role in sequential recommender systems. SR models usually handle separate interaction sequences in each domain. Then SR models transfer the knowledge obtained from a source domain to a target domain [86, 103]. It is because data distribution is different across domains. C2DSR [87] constructs a global item-to-item graph across two domains to learn transitions between two domains. DDGHM [104] constructs local graphs and a global graph to capture intra-domain sequential transitions and inter-domain sequential transitions, respectively. And, DDGHM leverages hybrid metric learning to alleviate data sparsity and improve recommendation performance.
@@ 5.4 Fusion of Different Features
In SR models, multiple features are usually used. Different features have different fusion methods. These fusion methods can mainly be summarized into three types. Some specific details are shown in Figure 8. The first type leverages simple methods to fuse item embeddings and feature embeddings together, for example, adding them together is to obtain the input of a backbone model. The fusion is shown as follows:
The second type fuses different features in a backbone model [25], for example, leveraging a decoupled features fusion mechanism in a self-attention mechanism. The fusion is shown as follows:
The third type of methods first fuses different feature embeddings together, and then passes the fused embeddings through a backbone model to learn feature representations. Meanwhile, item representations are learned by the other backbone model. Finally, the third method fuses item representations and feature representations together for prediction. The fusion is described as follows:

### v2

@@ 4.2.1 Recurrent Neural Networks
Compared with traditional SR models, RNNs-based SR models not only can capture short-term sequential dependencies, but also can capture long-term sequential dependencies and non-linear dynamics. In recent years, some RNNs-based SR models have been proposed. GRU4Rec [24] applies RNNs to capture long-term dependencies in session-based data. Based on GRU4Rec, GRU4Rec+ [30] proposes a novel negative item sampling and a novel ranking loss function. Some RNNs-based SR models are designed to capture each sequence independently, thus ignoring the global structure of all sequences. To solve this problem, KrNN-P [31] incorporates neighbor sequences into RNNs to capture global relationships while refining local ones. To consider the dependencies among users’ sequences, MrRNN [32] introduces manifold regularization into RNNs based on the multi-facets of collaborative filtering. However, these RNNs-based models have some shortcomings. For example, they cannot avoid vanishing gradients while handling ultra-long interaction sequences.
@@ 4.2.2 Convolutional Neural Networks
In recent years, several CNNs-based SR models have been proposed. DACNN [33] leverages CNNs to capture local sequential features while considering item sequences. Caser [34] leverages convolutional filters to capture both general preferences and sequential patterns. CosRec [35] leverages 2D convolutional neural networks to model complex interaction relationships. However, CNNs-based SR models have certain shortcomings. For example, they might have a large number of convolutional filters, which will result in excessive computational resource consumption.
@@ 4.2.3 Transformer
Transformer architecture has two types: single-direction Transformer (i.e., uni-Transformer) and double-direction Transformer (i.e., bi-Transformer). SASRec [1] is currently the most popular SR model, which leverages the uni-Transformer. SASRec predicts the next interacted items based on historical interacted items. BERT4Rec [36] leverages bi-Transformer, which predicts the next interacted items according to historical interacted items and future interacted items. ALBERT4Rec and DeBERTa4Rec [37] demonstrate superior performance and training efficiency compared to BERT4Rec [36]. ADT [38] learns disentangled diverse interests using a Transformer-based encoder and decoder. However, the large weights of attention mechanism might not be accurate. Additionally, position encoding and noise inputs might negatively impact the recommendation performance. Therefore, AC-TSR [39] leverages contrastive learning to achieve noise reduction. DT4IER [40] leverages decision Transformer to solve user retention challenges. Due to the uncertainty of users’ sequential behaviors, STOSA [41] leverages stochastic Gaussian distribution to embed items and introduces a novel Wasserstein self-attention module to capture item-item relationships. MFGAN [42] uses a Transformer-based generator to recommend the next possible items and multiple discriminators to evaluate generated sub-sequences. SSE-PT [43] introduces user embeddings to realize the personalization of Transformer. STRec [44] leverages sparse Transformer to focus on the most relevant interactions while predicting target items. Meanwhile, STRec [44] replaces the self-attention mechanism with a cross-attention mechanism. MoSeR [45] captures motifs hidden in behavior sequences to model micro-structure features. PHGR [46] leverages a novel hyperbolic inner product operator to enable global (all user-item interactions) and local (each user’s interactions) graph representation learning in Poincaré ball. CBiT [47] leverages contrastive learning and bi-Transformer to learn more fine-grained preferences. DCRec [48] leverages cross-view contrastive learning to capture correlations among different users’ interaction sequences. DFAR [49] leverages a dual-interest disentangling layer to disentangle positive and negative interests for learning users’ transition patterns better. DEPS [50] leverages both users’ and items’ interaction sequences as the inputs to two Transformers for learning preferences better. MrTransformer [51] leverages a preference separation module to learn both users’ common and unique preferences. CRR-Transformer [52] applies a pre-trained transformer model to an online RL algorithm. Due to the uncertainty of interactions, DT4SR [53] leverages an elliptical Gaussian distribution to describe items and adopts Wasserstein distance to measure the similarity among distributions. Transformers4Rec [54] applies Transformer-based architectures for recommendation, such as GPT-2, XLNET, and Transformer-XL. PAUP [55] captures both short- and long-term patterns through a progressive attention distribution mechanism. In this model, long behavior sequences are segmented into a series of sub-sequences using a down-sampling convolution module. SAMRec [56] enhances the recommendation performance of Transformer from the perspective of loss geometry.
@@ 4.2.4 Diffusion Model
In recent years, several SR models based on diffusion model have been proposed. DiffuASR [57] leverages a diffusion-based framework to generate high-quality interactions guaranteeing the similarity between generated and original sequences. DiffuRec [58] is the first work that applies the diffusion model to sequential recommendation. By corrupting and reconstructing target item embeddings, DiffuRec [58] injects uncertainty during the recommendation process. Diff4Rec [59] leverages the diffusion model to augment data by adopting a curriculum scheduling strategy, including interaction augmentation and objective augmentation. However, Diffusion-based SR models have some weaknesses. Compared with Transformer-based SR models, diffusion-based SR models require more computational resources and training time.
@@ 4.2.5 Graph Neural Networks
GNNs-based SR models can learn complex high-order interactions and capture collaborative information. Therefore, GNNs-based SR models have received widespread attention. AutoSeqRec [60] leverages an encoder and three decoders to reconstruct user-item interaction and item transition matrices. CD-GES-SASRec [61] leverages a causal graph to distinguish causal and non-causal transitions. DGSR [62] constructs a user-item interaction graph to capture personal preferences and an item-item transition graph to model transitional relationships between adjacent items. Meanwhile, DGSR [62] injects high-order connections into graphs. MSSG [63] transforms items within a sequence into a star graph. Meanwhile, MSSG introduces an additional internal node to capture global information. MSSG overcomes the over-smoothing issue and realizes a linear time complexity. However, GNNs-based SR models might not fully capture sequential patterns and could be computationally expensive to train.
@@ 4.2.6 Multi-Layer Perceptions
Compared with Transformer, leveraging MLPs to calculate is more efficient (i.e., a linear time complexity). In recent years, some MLPs-based SR models have been proposed. FMLP-Rec [64] incorporates learnable filters into MLPs architecture for reducing noise. TRIMLP [65] disables triangle neurons in MLPs to account for chronological order in users’ interaction sequences. MLM4Rec [66] learns users’ global and local preferences by incorporating convolution operators into MLPs. However, these MLPs-based SR models struggle to capture complex sequential patterns.
@@ 4.4 Hybrid Models
Some SR models demonstrate significant improvements in recommendation performance by combining different techniques. For example, combining Transformer and GNNs can achieve better performance. As shown in Table 2, each technique offers distinct advantages. Therefore, the hybrid models can combine the strengths of different techniques. The hybrid models can be roughly categorized into three classes (i.e., CNNs + Transformer, GNNs + Transformer and other combined types). Integrating various neural network architectures enables SR models to achieve enhanced recommendation performance. Four kinds of commonly used combined models are depicted in Figure 6. Two different techniques can be leveraged in parallel. For example, CNNs and Transformer can extract short-term dependencies and long-term dependencies from user interaction sequences, respectively. And then the two kinds of dependencies are fused together. Alternatively, the two techniques can be used in a cascaded manner. For example, GNNs can be denoted as a feature extractor for extracting features, and then the extracted features can be fed into Transformer.
@@ 4.4.1 Combination of CNNs and Transformer
Transformer and CNNs exhibit unique abilities to capture different types of dependencies. Transformer has a good ability to capture long-term dependencies. CNNs are effective at modeling short-term dependencies and multiple-faced preferences. As shown in part (a) of Figure 6, CNNs and Transformer can be combined in a parallel manner, which is formalized as follows:
The final preferences can be fed into a prediction module for recommending personalized items for user . Meanwhile, as shown in part (b) of Figure 6, CNNs and Transformer can be integrated in a cascaded manner, which is formalized as follows:
@@ 4.4.2 Combination of GNNs and Transformer
GNNs and Transformer excel at modeling graph structure and sequences, respectively. Transformer can capture temporal patterns and long-term dependencies. Meanwhile, GNNs can capture structural information and users’ long-term preferences. Therefore, combining GNNs and Transformer enables SR models to achieve better performance. As shown in part (d) of Figure 6, GNNs and Transformer can be combined in a parallel manner, which is formalized as follows:
In the meantime, as shown in part (c) of Figure 6, GNNs and Transformer can be integrated in a cascade manner, which is formalized as follows:
@@ 4.4.3 Other Combinations
There are also some SR models combining various other techniques. RANN [81] leverages the strengths of RNNs and the self-attention mechanism in Transformer to capture user preferences from different views. RCNN [82] combines the advantages of RNNs and CNNs, in which RNNs can capture long-term dependencies and CNNs can capture short-term sequential patterns. NISRec [83] captures and aggregates the long-term and short-term intentions of neighbor users. Then NISRec [83] fuses these intentions for recommendation. Mixture networks can leverage multiple experts to learn diverse interests. For example, M3 [84] leverages a mixture of experts, where each expert specializes in capturing a specific temporal range. These experts are dynamically combined through a learned gating mechanism. Therefore, M3 [84] can capture diverse behavior patterns effectively.
By combining different techniques, SR models can achieve better performance. However, these models have several drawbacks. Firstly, they are more complex and computationally expensive. Secondly, different techniques might capture conflicting preferences. For example, Transformer and GNNs capture distinct preferences, which may conflict. Thirdly, different techniques might converge at different rates.
@@ 5 SR with Side Information
In this section, our survey first discusses SR models combining IDs and one of the general features (i.e., categorical features and numerical features). Then, our survey shows SR models combining IDs and one of the graph-structure features (i.e., knowledge graph and social networks). Next, our survey introduces SR models combining IDs and reviews. Finally, our survey shows SR models combining IDs and multiple feature types.
@@
| MLPs-based SR models | Computational efficiency, a linear time and space complexity, simple model architecture | Limited performance on ultra-long sequences, are weak in capturing sequential dependencies |
@@
| TCNs-based SR models | Can preserve sequence order during convolution process and conduct efficient sequence modeling | Are difficult to capture long-term sequential dependencies, high complexity for hyper-parameter searching |
@@ 5.1 One Categorical or Numerical Feature
Many SR models combine either a categorical or a numerical feature with item IDs to recommend items for users. Some specific methods are shown in Figure 7. From Figure 7, we can know that these methods can be grouped into four main types. As shown in part (a) of Figure 7, the first method leverages two backbones to learn two kinds of preferences from one sequence containing item IDs and the other sequence containing feature , respectively. These preferences are then fused using methods such as contrastive learning, concatenation, addition, gate addition, or a cross-attention mechanism. The fusion process is formalized as follows:
As shown in part (b) of Figure 7, the second method leverages the value of feature to split the original sequence containing item IDs into sub-sequences. For example, some SR models split hybrid interaction sequences into separate interaction sequences in each domain [85, 86]. Different preferences are then extracted from these sub-sequences and fused together to predict personalized items. The fusion process is formalized below:
It is noted that the fusion usually exhibits a certain directionality, where user preferences are transferred from a source domain to a target domain [87]. As shown in part (c) of Figure 7, the third method first leverages simple techniques to combine item embeddings and feature embeddings. Then, the fused embeddings are fed to a backbone model for recommendation. For example, adding them together is to obtain the input of a backbone model. The fusion process is formalized as follows:
As shown in part (d) of Figure 7, the last method integrates item embeddings with feature embeddings in a backbone model. For example, leveraging feature embeddings changes the weights in attention matrices. The fusion process is formalized as follows:
Though all four kinds of methods can obtain final user ’s preferences, they vary in terms of complexity and efficiency [87, 88]. Meanwhile, as shown in part (e) of Figure 7, the predicted goal might contain the feature of the next interacted item [89, 25].
@@ 5.1.1 Behaviors
Different behaviors show users’ different preferences. Click behaviors usually reflect users’ short-term preferences. Favorite behaviors usually reflect users’ middle-term preferences. Purchase behaviors usually reflect users’ long-term preferences. Meanwhile, dependencies exist among these different behaviors. MB-STR [90] injects users’ behaviors into Transformer to capture multi-behavior heterogeneous dependencies. PBAT [91] models users’ personalized behavior patterns and behavioral collaborations by a personalized behavior pattern generator and a behavior-aware collaboration extractor, respectively. GHTID [92] leverages a global heterogeneous graph (constructed from all interaction sequences) and some local heterogeneous graphs (constructed from each interaction sequence) to explicitly learn heterogeneous item transitions.
@@ 5.1.4 Domain
Domain information plays a crucial role in sequential recommender systems. SR models usually first handle separate interaction sequences in each domain. Then SR models transfer the obtained knowledge from a source domain to a target domain [86, 103]. It is because data distribution is different across domains. C2DSR [87] constructs a global item-to-item graph across two domains to learn transitions between two domains. DDGHM [104] constructs local graphs and a global graph to capture intra-domain sequential transitions and inter-domain sequential transitions, respectively. And, DDGHM leverages hybrid metric learning to alleviate data sparsity and improve recommendation performance.
@@ 5.4 Fusion of Different Features
In SR models, multiple features are usually used to improve the recommendation performance. Different features can be fused by different fusion methods. These fusion methods can mainly be summarized into three types. Some specific details are shown in Figure 8. The first type leverages simple methods to fuse item embeddings and feature embeddings together, for example, adding them together is to obtain the input of a backbone model. The fusion is shown as follows:
As shown in part (b) of Figure 8, the second type fuses different features in a backbone model [25], for example, leveraging a decoupled features fusion mechanism in a self-attention mechanism. The fusion is shown as follows:
As shown in part (c) of Figure 8, the third type first fuses different feature embeddings together, and then passes the fused embeddings through a backbone model to learn feature representations. Meanwhile, item representations are learned by the other backbone model. Finally, the item representations and feature representations are fused together for prediction. The fusion is described as follows:
We will introduce common features fusion, context information fusion and special features fusion one by one.